Laurie invested some money at 12% and some more at 14%. She invested $50,000,000 more at 14% than shed did at 12%. Her total yearly interest from the two investments is $13,500,000. How much did she invest at each rate?
Answer by nerdybill(7384)

Laurie invested some money at 12% and some more at 14%. She invested $50,000,000 more at 14% than shed did at 12%. Her total yearly interest from the two investments is $13,500,000. How much did she invest at each rate?
.
Let x = amount ($) invested at 12%
then
x+50000000 = amount ($) invested at 14%
.
.12x + .14(x+50000000) = 13500000
.12x + .14x+7000000 = 13500000
.26x = 6500000
x = 6500000/.26
x = $25,000,000 (invested at 12%)
.
amount invested at 14%:
x+50000000 = 25000000+50000000 = $75,000,000